The typical dosage of statin drugs for the treatment of high cholesterol is 10 mg.

Assuming a total blood volume of 5.5 L , calculate the concentration of drug in the blood in units of % (w/v). Express your answer using two significant figures.
Answer #1

Answer:

The amount of drug given = 10mg or 0.01 g

Volume of blood = 5.5 L or 5500 mL

Now the %(w/v) is calculated/represented as (g of solute/100 mL of solvent) x 100

%(w/v) = (0.01/5500) x 100 = 1.8 x 10-4 g/100mL (%)  

The same can be written in other form also %(w/v) = 0.18 mg/100mL (%)  

Please let me know, if you have any doubt by commenting below the answer.

Thanks